# A Transformer Model for Predicting Chemical Reaction Products from Generic Templates

### v2

@@ IV-A Model and Implementation Details
The general architecture of ProPreT5 is illustrated in Figure 1. A T5 model was chosen for its ease of use, relatively lightweight architecture, and ability to handle large datasets even with modest training resources. The model is very similar to the original T5 [22], encompassing an encoder-decoder structure. The encoder and decoder blocks each consisting of 6 layers with a hidden size of 512. Each block includes self-attention and feed-forward layers, with 8 attention heads. Feed-forward layers have an intermediate size of 2048 and use ReLU activation. In addition to these layers, the decoder includes cross-attention layers and masked self-attention. Relative positional embeddings are also used to capture token relationships.
Both the encoder and decoder share a vocabulary embedding layer with a vocabulary size of 243 tokens, as both the input and output use the same notation—SMILES for reactants and product and either SMILES or SMARTS for reactions. The relatively small vocabulary size results from character-level tokenization, which enhances ProPreT5’s flexibility. This approach allows most datasets to be used without retraining the model, unless a new character is added—unlikely given that the vocabulary covers a substantial portion of known chemistry. Character-level tokenization was selected to address the challenge of ensuring pre-trained models have the correct tokens, as adding and finetuning new tokens is costly. Additionally, a larger Byte-Pair Encoding (BPE) tokenizer [32] with around 10,000 tokens was trained and tested but this led to generalization issues. This compact and efficient model enables the generation of reaction outcomes.
As demonstrated in Figure 1, the reactants and the reaction are distinguished with a separation token used to separate both the reactants from one another and the reaction from the reactants. Additionally, type embeddings are applied to each input, following the approach in [11]. These embeddings distinguish between different input and output types, making it easier to incorporate new entry types, such as reagents, without confusing the model. A separate trainable embedding layer is used to map these type tokens to embedding vectors. ProPreT5 can also make predictions in a template-free setting, meaning the reaction templates would be excluded from the encoder input sequence.
TABLE II: Comparison of Reaction Product Prediction Accuracy
Comparisons were not provided for the template-based versions, as none of the models included the necessary tokens for reaction templates. Additionally, the Molecule Transformer code was unavailable, and we were unable to run the Augmented Transformer. (*) Results were obtained without fine-tuning. Results with citations were directly taken from the corresponding papers.
Dataset Reaction Template Molecule Transformer Augmented Transformer Chemformer ProPreT5
USPTO MIT Mixed Template-based - - - 99.8%
USPTO MIT Mixed Template-free 88.6% [6] 90.0% [7] 90.9% [8] 87.2%
BRS Template-based - - - 85.8%
BRS Template-free - - 5.6%* 8.8%*
@@ IV-B Computational Resources and Training Setup
The training and evaluation of ProPreT5 were completed on a high-performance computing cluster equipped with NVIDIA V100 GPUs with 32 GB of memory, providing sufficient capacity to handle large datasets. ProPreT5 supports distributed training and was trained in parallel on 16 GPUs for 20 hours. Unlike some models in the literature, this setup was lightweight and time-efficient, yet even this relatively short 20-hour training provided satisfying results.
@@ V Results and Discussion
ProPreT5 was trained in both template-free and template-based settings. The template-free training primarily served as a sanity check, allowing us to compare ProPreT5’s performance with existing models in the literature. This step was essential to confirm that the model was functioning as expected and could be reliably used with BRS. The goal was not to improve the USPTO MIT baseline but to ensure the model’s performance aligned reasonably well with prior work, despite using an economical training approach with minimal data augmentation. Specifically, we augmented the dataset by a factor of four using non-canonical SMILES as proposed in [33], where multiple non-canonical forms of reactants were paired with the same product. Additionally, costly pretraining methods, commonly employed in the existing literature, were not used in this case.
@@
Table II presents the exact match accuracy of different models trained on various datasets. In the USPTO MIT Mixed dataset, each reactant-reaction combination corresponds to exactly one product, so if the model did not generate that specific product, the prediction was considered incorrect. In contrast, in the BRS dataset, a single reactant-reaction combination can lead to multiple possible products, similar to how multiple correct translations can exist for a machine translation task. Therefore, in the case of BRS, a generation was considered correct if the model predicted any one of the possible products.
V-A Template-Free Reaction Product Prediction
@@
As seen in Table II, after only 20 epochs of training with a relatively economical setup, ProPreT5 achieved an exact match accuracy of 87.2% on the template-free USPTO MIT Mixed baseline. This result is particularly impressive considering the hundreds of epochs required by the other three models. This accuracy was sufficient to confirm that the model was functioning as expected and ready to tackle the new, ambitious task at hand.
Predicting reaction products for the BRS dataset is inherently more complex than for the USPTO dataset. Without the use of reaction templates, accurately predicting products becomes even more challenging. Since a single molecule can act as a reactant in many different reactions, the model faces significant limitations in predicting the correct transformation. As a result, the model can only learn general transformation patterns and apply them randomly, which makes it highly unlikely that the generated product matches the expected outcome. The results obtained without reaction templates for the BRS dataset highlight the complexity of this task in the absence of templates.
To provide a point of comparison, a test epoch was run on Chemformer using the publicly available code from [8], without any fine-tuning. Similarly, ProPreT5 was also tested without fine-tuning. We would have liked to offer the same comparison with the other two models, but the code for the Molecule Transformer was not publicly available, and we were unable to successfully run the code for the Augmented Transformer.
The low percentage of correct predictions by both models underscores the difficulty of this task and suggests that in real-world scenarios where a reactant can be involved in multiple different reactions, models trained in a template-free setting struggle to generalize and produce accurate predictions.
(a) Mean Aggregation of Importance Matrices
(b) Max Aggregation of Importance Matrices
Figure 2: Interpretability Analysis of Template-Based Generations of ProPreT5 on USPTO and BRS
(a) Reactant Tokens
(b) Reactant Template Tokens
(c) Product Template Tokens
Figure 3: Importance of Each Token in Generation for a Single Example from USPTO
V-B Template-Based Reaction Product Prediction
When it comes to template-based prediction, ProPreT5 achieves a near-perfect accuracy of 99.8% on the USPTO MIT Mixed baseline, suggesting that the task becomes trivial since the templates are highly specific and essentially enumerate the atoms in the product. In contrast, for the BRS dataset, the templates provide essential information necessary for generating the correct product. Without this information, making accurate predictions becomes an inherently illogical task. Given that the same reactants can participate in multiple reactions within BRS and yield different products, the model must effectively interpret the information encoded in the generic reaction templates to make accurate predictions. Despite this complexity, ProPreT5 achieves an impressive accuracy of 85.8%, highlighting its effectiveness even in more difficult prediction scenarios.
To confirm the intuition about template-based predictions on USPTO and BRS, we focused on understanding how the model utilizes the information in the input sequence. For the analysis, we used Inseq [34], a tool for attributing importance scores to input tokens, which helps understand how the model generates its outputs. Integrated gradients were used as the attribution method. This analysis was conducted to enhance the interpretability of ProPreT5.
In Figure 2, the template-based generations were analyzed. The input sequence was divided into three subcategories: Reactants, the reactant template, and the product template of the reaction (typically separated by >>much-greater-than$>$ > in SMARTS notation). As a result, three importance score matrices were obtained using Inseq for each example, aligning with each of the three input sections. These matrices were either max-aggregated or mean-aggregated to provide a general overview of the entire test dataset. The violin plots in Figure 2 display the distribution of mean and max importance scores for the template-based predictions on the BRS and USPTO datasets.
As can be deduced by looking at both Figure 2(a) and Figure 2(b), most of the information in the input sequence is not utilized for generation on USPTO. Only a few tokens carry the essential information. These results support the earlier hypothesis that when generating a product, the model primarily focuses on a handful of tokens from the reactant part and the product template part of the input. The model focuses mainly on specific tokens, while the rest of the input contributes little to the generation.
Figure 2 shows the entire dataset, while Figure 3 highlights an example where darker colors indicate higher contribution to the output, and lighter colors indicate less contribution. This example from USPTO demonstrates that the contribution of the reactant pattern is minimal, as the target product is almost explicitly encoded in the reaction SMILES template. The model doesn’t need to find an alignment between the template reactants and the actual reactants to produce the result. This explains the near-perfect accuracy of 99.8% and why the task is trivial.
In contrast, the BRS dataset presents a more implicit relationship between the product and the reaction template. The mean aggregated importance graphs in Figure 2(a) for both parts of the reaction template are nearly identical, highlighting the model’s need to rely equally on both parts to make a correct prediction.
V-C Assesment Metrics for Template-Based ProPreT5
The template-based ProPreT5 model proved itself to be a good compromise between overly specific templates and template-free approaches. Its performance was further evaluated across various metrics, as presented in Table III.
Accuracy, as previously defined, refers to the exact match accuracy. For the BRS dataset, where multiple products can result from a single reactant-reaction combination, a prediction is considered correct if the model generates any one of the possible outcomes. Accuracy is then calculated as the ratio of correct predictions to the total number of predictions.
Validity evaluates whether the predicted molecules can be successfully transformed into molecular graphs using RDKit [26]. If a molecule violates the fundamental construction rules of organic chemistry, it cannot be parsed by RDKit and is thus considered invalid. Validity is determined as the ratio of parsable molecules to the total number of predictions.
On the other hand, RDKit isn’t capable of checking the realism of a product, and just because a molecule is valid doesn’t necessarily mean it’s realistic. Realism assesses whether the predicted molecules resemble known real-world molecules. This is achieved using subgraphs filters [29], which eliminate molecules containing groups or cycles never encountered in datasets of billions of real molecules. Realism is calculated as the proportion of predictions that pass these filters out of the total number of predictions.
The generic reactions from BRS are capable of generating valid but unrealistic products. As explained earlier in Section III-B, we filtered out those unrealistic products to construct the dataset. Table III demonstrates that ProPreT5, trained on this dataset, has the ability to generate valid, accurate, and realistic products, highlighting its utility in exploring and expanding the chemical space while respecting the rules of organic chemistry.
TABLE III: Metrics for Template-Based ProPreT5
Accuracy Validity Realism
85.8% 99.3% 91.0%
@@ VI Conclusion
In this study, we introduced the Broad Reaction Set (BRS), a novel generic reaction set designed to provide a more realistic and versatile benchmark. This important contribution addresses key limitations of the widely used USPTO dataset. Until now, template-based reaction prediction models relied on overly specific reaction templates, limiting their real-world applicability that we seek in cheminformatics.
The relevance of the BRS dataset was supported by the presented ProPreT5, a flexible reaction product prediction model capable of generating valid, realistic, and accurate products. Its flexibility comes not only from training on a dataset with generic reactions but also from its character-level tokenization, which enables easy adaptation to other datasets without costly retraining processes. ProPreT5 is a major contribution as it differs from existing models by predicting products in a way that can be applied to real-world chemical problems. We also demonstrated that obtaining satisfactory results in lighter training setups is possible.
Future work will move beyond single-step product predictions and explore multi-step organic synthesis routes while continuing to use a template-based approach with ProPreT5. Additionally, we will focus on improving the prediction accuracy to avoid propagating the error in a multi-step setup.
@@ Acknowledgment
The Acknowledgment section will be left empty at this time in order to maintain the anonymity of the submission.

### v3

@@ IV-A Model and Implementation Details
The general architecture of ProPreT5 is illustrated in Figure 1. A T5 model was chosen for its ease of use, relatively lightweight architecture, and ability to handle large datasets even with modest training resources. The model is very similar to the original T5 [21], encompassing an encoder-decoder structure. The encoder and decoder blocks each consist of 6 layers with a hidden size of 512. Each block includes self-attention and feed-forward layers, with 8 attention heads. Feed-forward layers have an intermediate size of 2048 and use ReLU activation. In addition to these layers, the decoder includes cross-attention layers and masked self-attention. Relative positional embeddings are also used to capture token relationships.
Both the encoder and decoder share a vocabulary embedding layer with a vocabulary size of 243 tokens, as both the input and output use the same notation. The relatively small vocabulary size results from character-level tokenization, which enhances ProPreT5’s flexibility. This approach allows most datasets to be used without retraining the model, unless a new character is added, unlikely given that the vocabulary covers a substantial portion of known chemistry. Character-level tokenization was selected to address the challenge of ensuring pretrained models have the correct tokens, as adding and fine-tuning new tokens is costly.
As demonstrated in Figure 1, the reactants and the reaction are distinguished with a separation token used to separate both the reactants from one another and the reaction from the reactants. Additionally, type embeddings are applied to each input, following the approach in [10]. These embeddings distinguish between different input and output types, making it easier to incorporate new entry types, such as reagents or environmental conditions, without confusing the model. A separate trainable embedding layer is used to map these type tokens to embedding vectors. ProPreT5 is trained separately to make predictions in both template-free and template-based settings: in the template-free setting, reaction templates are excluded from the encoder input; in the template-based setting, the model has access to the reaction template as part of the input.
@@ IV-B Computational Resources and Training Setup
The training and evaluation of ProPreT5 were completed on a high-performance computing cluster equipped with NVIDIA V100 GPUs with 32 GB of memory, providing sufficient capacity to handle large datasets. ProPreT5 supports distributed training and was trained in parallel on 4 nodes, each equipped with 4 GPUs, for a total of 6 hours. Unlike other models that require extensive computational resources and longer training times, this setup was relatively lightweight and time-efficient. Despite the short training duration, ProPreT5 achieved competitive performance.
@@ V Results and Discussion
ProPreT5 was trained in both template-free and template-based settings. While the USPTO MIT Mixed benchmark is almost exclusively addressed in the template-free setting in the current literature, our use of this setting served primarily as a sanity check. We applied template-free training on USPTO MIT Mixed not to improve upon existing baselines, but to verify that ProPreT5 behaved as expected and achieved performance reasonably aligned with prior work. This step was essential to confirm that the model was functioning correctly before applying it to the more challenging and novel dataset, which is the primary focus of this study.
V-A Product Prediction on USPTO MIT
In the template-free setting using the USPTO MIT Mixed dataset, we followed the standard approach in the literature where product prediction is performed without access to a reaction template. To improve generalization, we augmented the input molecules by a factor of four using non-canonical SMILES, as proposed in [32]; multiple non-canonical forms of each reactant were paired with the same product. This data augmentation is often employed in the existing literature [6, 7, 8]. On the other hand, unlike some of the existing studies that rely on costly pretraining strategies, our model was trained from scratch, directly on USPTO MIT Mixed.
@@ V-A Product Prediction on USPTO MIT
Table II presents the exact match accuracy of different models trained on the USPTO MIT Mixed dataset. For this dataset, each reactant-reaction combination corresponds to exactly one product, so if the model did not generate that specific product, the prediction was considered incorrect.
@@
As shown in Table II, after only a few hours of training on a relatively economical setup, ProPreT5 achieved an exact match accuracy of 87.9% in the template-free setting on the USPTO MIT Mixed benchmark. This result is particularly notable given that the other three models required extensive training over hundreds of epochs. The accuracy was sufficient to confirm that the model was functioning as expected and ready to be applied to the more ambitious BRS dataset.
We also explored the template-based setting for the same benchmark. However, the other models could not be evaluated in this setting, as they lacked support for reaction template tokens. Introducing new tokens and retraining those models would have required significant effort and computational resources. However, ProPreT5 achieved a near-perfect accuracy of 99.8% in the template-based setting on USPTO MIT Mixed, supporting our earlier claim that the templates in this benchmark are highly specific and deterministic, effectively enumerating the atoms in the product.
TABLE II: Prediction Accuracy on USPTO MIT Mixed
Model Template-based Template-free
Molecule Transformer - 88.6% [6]
Augmented Transformer - 90.0% [7]
Chemformer - 90.9% [8]
ProPreT5 99.8% 87.9%
Note: Template-based comparisons were omitted because none of the baseline models natively support reaction templates, and extending them with new tokens and retraining poses non-trivial challenges. Results with citations were directly taken from the corresponding papers.
V-B Product Prediction on BRS
Predicting reaction products for the dataset generated using the reactions from BRS is inherently more complex than for the USPTO MIT Mixed dataset. Unlike the highly specific templates in USPTO, the BRS templates are generic and do not explicitly enumerate the input or output molecules. Instead, they describe general transformation mechanisms that must be learned and generalized across diverse molecular contexts. The model must not only understand these transformations but also recognize that they can apply to a wide variety of molecules and to different regions within a single molecule. By design, this makes the task significantly more challenging. Using BRS templates, a single reactant-reaction combination can lead to multiple possible products, similar to how multiple correct translations can exist for a machine translation task. Therefore, in the case of BRS, a generation is considered correct if the model predicted any one of the possible products. It is important to note that, in this setting, no training sample shares the same reactant–reaction input as any test sample, a condition we carefully ensured.
The difficulty of the task increases further in the absence of reaction templates. Since a single molecule can serve as a reactant in many distinct reactions, the model lacks critical contextual cues and must rely on general transformation patterns. As a result, predictions become less reliable: the model may propose a chemically plausible product, but one that does not correspond to the specific reaction intended to produce the target product. This limitation is evident in the template-free results on the BRS-Base dataset shown in Table III, where both models achieve prediction accuracies below 70%.
For the template-free comparison, we selected Chemformer [8], as its code and trained weights were publicly available and it achieved the highest reported Top-1 accuracy on the USPTO MIT Mixed benchmark. To evaluate its performance on our benchmark, we fine-tuned Chemformer without reaction templates on the BRS-Base training set. Using BRS-Aug training set in the template-free setting is unnecessary, as the only difference from BRS-Base lies in the reaction templates, which are excluded in this configuration. The underwhelming accuracy by both models in the template-free setting highlights a key limitation: in real-world scenarios where a single reactant can participate in a wide range of reactions, models trained without templates may lack the reliability needed for accurate prediction.
TABLE III: Prediction Accuracy on BRS
Training Reaction Template Chemformer [8] ProPreT5
BRS-Base Template-free 69.2% 68.8%
BRS-Base Template-based - 91.2%
BRS-Aug Template-based - 95.1%
Note: Comparisons were not provided for the template-based versions, as Chemformer [8] does not support reaction templates by default, and adding new tokens followed by retraining the model would be prohibitively expensive. The template-based result on BRS-Base was obtained by fine-tuning Chemformer on top of its existing training on the USPTO MIT Mixed dataset. The reported accuracies are based on the BRS-Base test set, while template augmentation (BRS-Aug) is applied only during training to improve prediction performance.
When it comes to template-based prediction, which is the main challenge addressed in this paper, we aim to propose a language model capable of handling and understanding the SMARTS notation. This capability is essential not only for accurate product prediction but also for enabling a variety of downstream tasks in the future. To evaluate this, we first trained the model on the non-augmented version of the dataset BRS-Base. The model achieved a solid accuracy of 91.2% on the test set. However, given the limited number of reaction templates, there is a risk that the model simply memorizes the reactions rather than learning the underlying transformation rules encoded in the SMARTS notation.
To address this concern, we trained the model on the augmented version of the training set BRS-Aug. This version includes multiple syntactic variants of each reaction template as explained in detail in Section III-C, providing greater diversity and encouraging the model to generalize. As a result, the model’s accuracy improved to 95.1% on the same test set, indicating that the augmentation strategy indeed helps the model capture a more nuanced understanding of the reaction rules.
This effect becomes particularly evident when we assess the model’s robustness by removing a specific reaction template from the training set, specifically, Reaction 5 from Table I, along with all its augmented variants. As shown in Table IV, the model trained on the non-augmented BRS-Base dataset without the Reaction 5 suffers a significant drop in performance when evaluated on that same reaction, suggesting that it had overfitted to the limited examples without truly learning the SMARTS syntax. In contrast, the model trained with augmented templates maintains much better performance, providing strong evidence that the augmentation strategy enhances generalization and improves the model’s ability to learn the SMARTS language in a meaningful way.
TABLE IV: Prediction Accuracy of ProPreT5 on Reaction 5 Removed from Training Set
Training Subset Accuracy on Reaction 5
BRS-Base without Reaction 5 23.3%
BRS-Aug without Reaction 5 54.7%
Note: Reaction 5 and all its augmented variants were removed from the training sets. The table reports the model’s accuracy on test samples involving Reaction 5.
@@ VI Conclusion
In this study, we introduced the Broad Reaction Set (BRS), a novel collection of generic reaction templates designed to provide a more realistic and versatile benchmark for reaction prediction. This contribution addresses key limitations of widely used datasets such as USPTO, which rely on overly specific reaction templates and therefore limit real-world applicability, an essential requirement in cheminformatics.
We also proposed a novel augmentation strategy for the SMARTS notation, which, to the best of our knowledge, has not been explored previously. This strategy was applied to the BRS templates, resulting in two training sets: BRS-Base, using the original templates, and BRS-Aug, incorporating augmented variants. While developed for the 20 generic reactions presented in this study, the augmentation strategy is generalizable and can be applied to other SMARTS-based reaction templates.
To evaluate the impact of the proposed datasets, we introduced ProPreT5, a flexible T5-based model for reaction product prediction. Its flexibility stems not only from being trained on generic reactions but also from its character-level tokenization, which enables seamless adaptation to other datasets without requiring costly retraining. ProPreT5 is a major contribution, as it is the first language model capable of directly handling and applying the SMARTS notation. This opens the door to new applications: thanks to its strong predictive performance, ProPreT5 can serve as a faster and more scalable alternative to RDKit for product prediction, particularly in scenarios where batching is required, something that RDKit’s template application does not support efficiently. Although its accuracy is not yet perfect, ProPreT5 offers a compelling balance of speed and flexibility. Furthermore, it provides a strong baseline for future downstream tasks involving SMARTS-based reasoning.
It is important to emphasize the broader utility of developing models capable of interpreting chemical reactions from generic templates. By learning the underlying rules of chemistry, such models are better prepared for downstream tasks that require a deeper understanding of reaction mechanisms. Moreover, unlike rule-based tools such as RDKit, which often fail when faced with noisy or imperfect input data, these models offer greater flexibility and generalizability. Their ability to handle variability and incomplete information makes them valuable assets for real-world applications where input quality cannot always be guaranteed.
Finally, we showed that satisfactory results can be obtained with lightweight training configurations, avoiding the need for extensive pretraining. In future work, we plan to extend this approach to multi-step synthesis planning, leveraging the template-based framework of ProPreT5. We also aim to further improve prediction accuracy in order to minimize error propagation in multi-step workflows.
@@ Acknowledgment
This work was supported by the University of Angers, and the French Ministry of Education and Research (JL PhD grant). This work was performed using HPC resources from GENCI-IDRIS (Grant 2024-AD011014840R1)